# A Bayesian workflow for securitizing casualty insurance risk

### v2

@@ 4 Discussion
More generally, this is the first paper of its kind to introduce a fully worked-out Bayesian workflow to address casualty insurance loss modeling. In the same way that models used throughout actuarial science have been studied, extended, and critiqued, we believe that actuarial workflows should be formalized and rigorously studied in an effort to both improve and disseminate them. To support future research in this vein, we have made all of the scripts used throughout this paper available in a repository (https://github.com/LedgerInvesting/bayesian-workflow-paper-2024). In doing this work, it has also become clear to us that future research on ILS (and throughout actuarial science more generally) would greatly benefit from the availability of richer open datasets that can be used to test model assumptions. Standardized open datasets would allow for researchers across institutions to compare models against common benchmarks, accelerating model developent throughout actuarial science. The Meyers (2015) dataset is a great starting place, but any results based on it should be taken with a grain of salt given its size and lack of recency. Larger datasets with longer, more recent histories and more lines of business are needed to ensure that results are generalizable.

### v3

@@ 5 Discussion
More generally, this is the first paper of its kind to introduce a fully worked-out Bayesian workflow to address casualty insurance loss modeling. In the same way that models used throughout actuarial science have been studied, extended, and critiqued, we believe that actuarial workflows should be formalized and rigorously studied in an effort to both improve and disseminate them. To support future research in this vein, we have made all of the scripts used throughout this paper available in a repository (https://github.com/LedgerInvesting/bayesian-workflow-paper-2024). In doing this work, it has also become clear to us that future research on ILS (and throughout actuarial science more generally) would greatly benefit from the availability of richer open datasets that can be used to test model assumptions. Standardized open datasets would allow for researchers across institutions to compare models against common benchmarks, accelerating model developent throughout actuarial science. The Meyers (2015) dataset is a great starting place, but any results based on it should be taken with a grain of salt given its size and lack of recency. Similarly, the Swiss liability triangle from Balona and Richman (2020) is a good candidate for testing model performance given its 18 year history. However, it is just a single triangle, and the most recent accident years available are from 2015. Larger datasets with longer, more recent histories and more lines of business are needed to ensure that published backtest results are generalizable. Although many private organizations have access to such data (e.g., through purchase agreements with data aggregators), licensing restrictions prohibit open use of these datasets. The impetus then is on the research community to advocate for more data sharing and collaboration, which could ultimately lead to more robust and reliable models in the field.